# Information-Theoretic Limits of Bistatic Integrated Sensing and Communication

### v3

@@ III-B Lower bounds of C(D)
In the bistatic ISAC system shown in Fig. 2, the estimator does not have access to the sent messages, resulting in larger estimation errors compared to the monostatic ISAC system, where the estimator knows the sent messages. To mitigate this, we leverage communication to assist sensing, meaning that the decoder at the SenRx can decode the sent messages to a certain extent. Based on the degree to which communication assists sensing at the SenRx, we propose three DnE strategies: blind estimation, partial-decoding-based estimation, and full-decoding-based estimation. Consequently, these strategies lead to three corresponding lower bounds.
@@
where the joint distribution of $SXYZ$ is given by $P_{X}P_{S}P_{YZ|XS}$ and the estimator $\hat{s}^{*}(z)=\arg\min_{s^{\prime}\in\hat{\mathbf{S}}}\sum_{s\in\mathbf{S}}P_{S|Z}(s|z)d\left(s,s^{\prime}\right)$.
@@
From the proof of Theorem 1 we can obtain $C(D)=\lim_{k\rightarrow\infty}C_{k}(D)=\sup_{k}C_{k}(D)\geq C_{1}(D)=\max_{P_{X}}\big{\{}I(X;Y|S)\big{|}\mathbb{E}\left[d\big{(}S,\hat{s}^{*}(Z)\big{)}\right]\leq D\big{\}}$, which completes the proof. ∎
@@
where $\hat{s}^{*}(u,z)=\arg\min_{s^{\prime}\in\mathbf{S}}\sum_{s\in\mathbf{S}}P_{S|UZ}(s|u,z)d\left(s,s^{\prime}\right)$. The joint distribution of $SUXYZ$ is given by $P_{UX}P_{S}P_{YZ|XS}$ for some pmf $P_{UX}$, and the cardinality of the auxiliary random variable $U$ satisfies $\left|\mathcal{U}\right|\leq|\mathcal{X}|+1$.
@@
Assume, without loss of generality, that $\left(W_{0},W_{1}\right)=(1,1)$ is sent. The average probability of error is $P_{1e}^{(n)}=\mathrm{P}\big{\{}(\hat{W}_{01},\hat{W}_{1})\neq\left(1,1\right)\text{ or }\hat{W}_{02}\neq 1\big{\}}$. Applying the jointly typical decoding at the decoder and following similar arguments as superposition coding inner bound, we can obtain that $P_{1e}^{(n)}$ tends to zero if $R_{0}<I\left(U;Z\right)-\delta(\epsilon)$, $R_{1}<I\left(X;Y|U,S\right)-\delta(\epsilon)$, and $R_{0}+R_{1}<I\left(X;Y|S\right)-\delta(\epsilon)$ hold, where $\delta(\epsilon)\rightarrow 0$ when $n\rightarrow\infty$.
@@
Note that for $SUXYZ$ distributed as $P_{UX}P_{S}P_{YZ|XS}$, $\hat{s}^{*}(u,z)$ is the optimal one-shot state estimator, where the term ‘one-shot’ means that the estimator estimates the state $s_{i}$ in time slot $i$ only as a function of $u_{i}$ and $z_{i}$. In addition, $\hat{s}^{*}(u,z)$ can be interpreted as a minimizer of the penalty function on the distortion measure $d$ and $P_{S|UZ}(s|u,z)$ is the posterior probability of $S$ when $(U,Z)$ is known. In particular, when the distortion measure $d$ is the Hamming distance, $\hat{s}^{*}(u,z)$ is a maximum a posteriori probability estimate. The proof on the optimality is given in Appendix B.
@@
where the joint distribution of $SXYZ$ is given by $P_{X}P_{S}P_{YZ|XS}$ and the optimal estimator $\hat{s}^{*}(x,z)=\arg\min_{s^{\prime}\in\hat{\mathbf{S}}}\sum_{s\in\mathbf{S}}P_{S|XZ}(s|x,z)d\left(s,s^{\prime}\right)$.
@@
Similar to the proof of Theorem 3, the conclusion is obtained by letting the SenRx decode the sent information $X$ and applying the optimal estimator $\hat{S}^{n}=\big{(}\hat{s}^{*}(\hat{x}_{1},z_{1}),\hat{s}^{*}(\hat{x}_{2},z_{2}),\ldots,\hat{s}^{*}(\hat{x}_{n},z_{n})\big{)}$. ∎
@@
Note that we adopt the average expected distortion defined in (1) as the distortion metric, which may not be more meaningful than the maximum distortion in some scenarios. However, they are equivalent in the case under consideration, i.e., all the single-letter bounds presented in the paper also apply to the maximum distortion measure. As is well known, the converse proof for the average metric can be directly extended to the case of the maximum metric. Therefore, in the following, we primarily analyze the lower bound based on the average metric. Specifically, based on the model in Fig. 2, we define that a maximum expected distortion as
$\Delta_{n}^{\max}\triangleq\max_{w}\Delta_{n}(x^{n}(w))=\max_{w}\frac{1}{n}\mathbb{E}\big{(}d\big{(}S^{n},\hat{S}^{n}(Z^{n})\big{)}|W=w\big{)}.$
@@
$\Delta_{n}(x^{n})=$ $\frac{1}{n}\mathbb{E}\big{(}d\big{(}S^{n},\hat{S}^{n}(Z^{n})\big{)}|W=w\big{)}\stackrel{{\scriptstyle(a)}}{{=}}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[d\big{(}S_{i},\hat{s}_{i}({Z_{i}})\big{)}\big{|}X_{i}=x_{i}\right]$
@@
$=$ $\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}d\big{(}S,\hat{s}(Z)|X_{i}=x_{i}\big{)}=\frac{1}{n}\sum_{i=1}^{n}\sum_{a\in\mathcal{X}}\mathbbm{1}[x_{i}=a]\mathbb{E}\big{[}d\big{(}S,\hat{{s}}(Z)\big{)}\big{|}X_{i}=a\big{]}$
@@ III-C Upper bounds of C(D)
An upper bound can be obtained by introducing a genie to the system that can feed the information sent to the estimator, as shown in Theorem 4 below.
@@
Note that when the SenRx is statistically stronger than the ComRx, i.e., mathematically, $X-Z-(Y,S)$ forms a Markov chain, we get that if the ComRx can decode the message sent, i.e., $R\leq I(X;Y|S)$, then the SenRx also can decode the message sent because of $R\leq I(X;Y|S)\leq I(X;Z)$. In this case, the estimator obtains the sent message and thus the capacity-distortion function $C(D)=C_{o}(D)$.
@@
$+I(U,V;Z)\big{|}\mathbb{E}\big{[}d(S,\hat{S}^{*}(U,V,Z))\leq D\big{]}\big{\}},$ (16)
where the joint distribution of $UVSXYZ$ is given by $P_{UV}P_{X|UV}P_{S}P_{YZ|XS}$, the estimator $\hat{s}^{*}(u,v,z)=\arg\min_{s^{\prime}\in\mathbf{S}}\sum_{s\in\mathbf{S}}P_{S|UVZ}(s|u,v,z)d\left(s,s^{\prime}\right)$, and the cardinalities of the auxiliary random variables $U$ and $V$ satisfy $\left|\mathcal{U}\right|\leq|\mathcal{X}|+2$ and $\left|\mathcal{V}\right|\leq|\mathcal{U}||\mathcal{X}|+1$, respectively.
@@
In the following, we consider the distortion constraint. For any estimator satisfying $D\geq\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\big{[}d\big{(}S_{i},\hat{S}_{i}(Z^{n})\big{)}\big{]}$, we have
$D\geq\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\big{[}d\big{(}S_{i},\hat{S}_{i}(U_{i},V_{i},Z^{n})\big{)}\big{]}\stackrel{{\scriptstyle(a)}}{{=}}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\big{[}d\big{(}S_{i},\hat{S}_{i}(U_{i},V_{i},Z_{i})\big{)}\big{]}\geq\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\big{[}d\big{(}S_{i},\hat{s}^{*}(U_{i},V_{i},Z_{i})\big{)}\big{]},$
where $(a)$ follows by the composition of the auxiliary random variable identification $U_{i},V_{i}$. Recall that the definition of the estimator in Theorem 3, $\hat{s}^{*}\big{(}(U_{i},V_{i},i),Z_{i}\big{)}=\hat{s}^{*}\left(U_{i},V_{i},Z_{i}\right)$ for $i=1,\cdots,n$. Using the law of total expectation and the definitions of $Q,S,U,V,Z$ above, we have
@@
$=\mathbb{E}\bigg{[}\mathbb{E}\Big{[}d\Big{(}S_{Q},\hat{s}^{*}\big{(}(U_{Q},V_{Q},Q),Z_{Q}\big{)}\Big{)}\Big{]}\Big{|}Q\bigg{]}$
$=\mathbb{E}\big{[}d\big{(}S,\hat{s}^{*}(U,V,Z)\big{)}\big{]},$ (17)
@@
Note that when the ComRx is statistically stronger than the SenRx, we have the following result.
Corollary 3.
@@
Since $(X,S)-Y-Z$ forms a Markov chain, we have $I\left(U;Y|S\right)\geq I\left(U;Y\right)\geq I\left(U;Z\right)$ holds for all $p(u,x)$. Hence, $I\left(U;Z\right)+I\left(X;Y|U,S\right)\leq I\left(U;Y|S\right)+I\left(X;Y|U,S\right)=I\left(X;Y|S\right)$ and $I(X;Y,Z|U,V,S)+I(U,V;Z)=I(X;Y|U,V,S)+I(U,V;Z)=I(X;Y|\tilde{U},S)+I(\tilde{U};Z)$. Also, we have $\mathbb{E}\big{[}d(S,\hat{S}^{*}(U,V,Z))=\mathbb{E}\big{[}d(S,\hat{S}^{*}(\tilde{U},Z))$. Furthermore, we get that the upper bound 2 coincides with the lower bound in Theorem 3. Thus, we obtain the capacity-distortion function $C(D)$, as shown in (18). ∎
@@
Corollary 3 implies that the optimality of the partial-decoding-based estimation strategy achieved by superposition coding in the degraded case. In addition, the result in Corollary 3 also holds when the Markov chain $X-(Y,S)-Z$ holds.
Referring to the lossy source coding for this ISAC system, we can get the following upper bound.
@@
Furthermore, similar to the converse proof in lossy source coding, we have $\sum_{i=1}^{n}I(S_{i};\hat{S}_{i})\leq I(S^{n};\hat{S}^{n})$. And since the Markov chain $S^{n}-Z^{n}-\hat{S}^{n}$ and the composition of auxiliary random variable $U_{i}$, we get $\sum_{i=1}^{n}I(S_{i};\hat{S}_{i})\leq I(S^{n};\hat{S}^{n})\leq I(S^{n};Z^{n})\leq\sum_{i=1}^{n}I(S_{i},U_{i};Z_{i})$, which completes the proof. ∎

### v4

@@ III-B Lower bounds of C(D)
In the bistatic ISAC system shown in Fig. 2, the estimator does not have access to the sent messages, resulting in larger estimation errors compared to the monostatic ISAC system, where the estimator knows the sent messages. To mitigate this situation, we leverage communication to assist sensing, meaning that the decoder at the SenRx decodes the sent messages to some extent, thereby facilitating the estimation based on the decoded information. Based on the degree to which communication assists sensing at the SenRx, we propose three decoding-and-estimation (DnE) strategies: blind estimation, partial-decoding-based estimation, and full-decoding-based estimation, which lead to three lower bounds on capacity-distortion function.
@@
where the joint distribution of $SXYZ$ is given by $P_{X}P_{S}P_{YZ|XS}$ and the estimator $\hat{s}^{*}(z)=\arg\min_{s^{\prime}\in\mathcal{S}}\sum_{s\in\mathcal{S}}P_{S|Z}(s|z)d\left(s,s^{\prime}\right)$.
@@
From the proof of Theorem 1 we obtain $C(D)=\lim_{k\rightarrow\infty}C_{k}(D)=\sup_{k}C_{k}(D)\geq C_{1}(D)=\max_{P_{X}}\big{\{}I(X;Y|S)\big{|}\mathbb{E}\left[d\big{(}S,\hat{s}^{*}(Z)\big{)}\right]\leq D\big{\}}$, which completes the proof. ∎
@@
where $\hat{s}^{*}(u,z)=\arg\min_{s^{\prime}\in\mathcal{S}}\sum_{s\in\mathcal{S}}P_{S|UZ}(s|u,z)d\left(s,s^{\prime}\right)$. The joint distribution of $SUXYZ$ is given by $P_{UX}P_{S}P_{YZ|XS}$ for some pmf $P_{UX}$, and the cardinality of the auxiliary random variable $U$ satisfies $\left|\mathcal{U}\right|\leq|\mathcal{X}|+1$.
@@
Assume, without loss of generality, that $\left(W_{0},W_{1}\right)=(1,1)$ is sent. The average probability of error is $P_{1e}^{(n)}=\mathrm{P}\big{\{}(\hat{W}_{01},\hat{W}_{1})\neq\left(1,1\right)\text{ or }\hat{W}_{02}\neq 1\big{\}}$. Applying the jointly typical decoding at the decoder and following similar arguments as superposition coding inner bound, we obtain that $P_{1e}^{(n)}$ tends to zero if $R_{0}<I\left(U;Z\right)-\delta(\epsilon)$, $R_{1}<I\left(X;Y|U,S\right)-\delta(\epsilon)$, and $R_{0}+R_{1}<I\left(X;Y|S\right)-\delta(\epsilon)$ hold, where $\delta(\epsilon)\rightarrow 0$ when $n\rightarrow\infty$.
@@
Note that for $SUXYZ$ distributed as $P_{UX}P_{S}P_{YZ|XS}$, $\hat{s}^{*}(u,z)$ is the optimal one-shot state estimator, which means that the estimator estimates the state $s_{i}$ in time slot $i$ only as a function of $u_{i}$ and $z_{i}$. In addition, $\hat{s}^{*}(u,z)$ can be interpreted as a minimizer of the penalty function on the distortion measure $d$ and $P_{S|UZ}(s|u,z)$ is the posterior probability of $S$ when $(U,Z)$ is known. In particular, when the distortion measure $d$ is the Hamming distance, $\hat{s}^{*}(u,z)$ is a maximum a posteriori probability estimate. The proof on the optimality is given in Appendix B.
@@
Similar to the proof of Theorem 3, the conclusion is obtained by letting the SenRx decode the sent information $X$ and applying the optimal estimator $\hat{S}^{n}=\big{(}\hat{s}^{*}(\hat{x}_{1},z_{1}),\hat{s}^{*}(\hat{x}_{2},z_{2}),\ldots,\hat{s}^{*}(\hat{x}_{n},\\ z_{n})\big{)}$. ∎
@@
Note that we adopt the average expected distortion defined in (1) as the distortion metric, which may not be more meaningful than the maximum distortion in some scenarios. However, they are equivalent in the case under consideration, i.e., all the single-letter bounds presented in the paper also apply to the maximum distortion measure. For converse statements, the maximum is more restrictive than the average. If the converse holds for the average, i.e., no code exists with a certain average performance in terms of error probability and distortion, then also no code exists with the maximum performance, which means the converse holds for the maximum. Therefore, in the following, we primarily analyze the lower bound (achievability) based on the average metric. Specifically, based on the model in Fig. 2, we define that a maximum expected distortion as
$\Delta_{n}^{\max}\triangleq\max_{w}\Delta_{n}\big{(}x^{n}(w)\big{)}=\max_{w}\frac{1}{n}\mathbb{E}\big{[}d\big{(}S^{n},\hat{S}^{n}(Z^{n})\big{)}|W=w\big{]}.$
@@
$\Delta_{n}(x^{n})=\frac{1}{n}\mathbb{E}\big{[}d\big{(}S^{n},\hat{S}^{n}(Z^{n})\big{)}|W=w\big{]}\stackrel{{\scriptstyle(a)}}{{=}}\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\big{[}d\big{(}S_{i},\hat{S}_{i}({Z_{i}})\big{)}\big{|}X_{i}=x_{i}\big{]}$
@@
$=$ $\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\big{[}d\big{(}S,\hat{S}(Z)\big{)}|X_{i}=x_{i}\big{]}=\frac{1}{n}\sum_{i=1}^{n}\sum_{a\in\mathcal{X}}\mathbbm{1}[x_{i}=a]\mathbb{E}\big{[}d\big{(}S,\hat{{S}}(Z)\big{)}\big{|}X_{i}=a\big{]}$
@@ III-C Upper bounds of C(D)
An upper bound can be obtained by introducing a genie to the system that feeds the sent message to the estimator, as shown in Theorem 4 below.
@@
Note that when the SenRx is statistically stronger than the ComRx, i.e., the channel to the ComRx is physically degraded or stochastically degraded with respect to the channel to the SenRx, we get that if the ComRx can decode the message sent, i.e., $R\leq I(X;Y|S)$, then the SenRx also can decode the message sent because of $R\leq I(X;Y|S)\leq I(X;Z)$. In this case, the estimator obtains the sent message and thus the corresponding capacity-distortion function $C(D)=C_{o}(D)$.
In the following, we derive a new upper bound by introducing a genie-aided estimator, where the genie supplies the estimator with some prior information instead of the sent message.
@@
$+I(U,V;Z)\big{|}\mathbb{E}\big{[}d\big{(}S,\hat{S}^{*}(U,V,Z)\big{)}\big{]}\leq D\big{\}},$ (16)
where the joint distribution of $UVSXYZ$ is given by $P_{UV}P_{X|UV}P_{S}P_{YZ|XS}$, the estimator $\hat{s}^{*}(u,v,\\ z)=\arg\min_{s^{\prime}\in\mathcal{S}}\sum_{s\in\mathcal{S}}P_{S|UVZ}(s|u,v,z)d\left(s,s^{\prime}\right)$, and the cardinalities of the auxiliary random variables $U$ and $V$ satisfy $\left|\mathcal{U}\right|\leq|\mathcal{X}|+2$ and $\left|\mathcal{V}\right|\leq|\mathcal{U}||\mathcal{X}|+1$, respectively.
@@
In the following, we consider the distortion constraint. For each estimator satisfying $D\geq\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\big{[}d\big{(}S_{i},\hat{S}_{i}(Z^{n})\big{)}\big{]}$, we define a genie-aided estimator ${\hat{S}}^{\prime}_{i}(U_{i},V_{i},Z^{n})$, whose corresponding distortion is less than that of estimator ${\hat{S}}_{i}(Z^{n})$. Such an estimator exists since more information is exploited. Then, for each genie-aided estimator ${\hat{S}}^{\prime}_{i}(U_{i},V_{i},Z^{n})$, we construct a new genie-aided estimator ${\hat{S}}^{\prime\prime}_{i}(U_{i},V_{i},Z_{i})$, which satisfies ${\hat{S}}^{\prime}_{i}(U_{i},V_{i},Z^{n})={\hat{S}}^{\prime\prime}_{i}(U_{i},V_{i},Z_{i})$ due to the composition of the auxiliary random variable identification $U_{i},V_{i}$. Thus, we have
$D$ $\geq\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\big{[}d\big{(}S_{i},\hat{S}_{i}(Z^{n})\big{)}\big{]}$
$\geq\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\big{[}d\big{(}S_{i},{\hat{S}}^{\prime}_{i}(U_{i},V_{i},Z^{n})\big{)}\big{]}$
$=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\big{[}d\big{(}S_{i},{\hat{S}}^{\prime\prime}_{i}(U_{i},V_{i},Z_{i})\big{)}\big{]}$
$\geq\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\big{[}d\big{(}S_{i},\hat{s}^{*}(U_{i},V_{i},Z_{i})\big{)}\big{]}.$
Recall that the definition of the estimator in Theorem 3, $\hat{s}^{*}\big{(}(U_{i},V_{i},i),Z_{i}\big{)}=\hat{s}^{*}\left(U_{i},V_{i},Z_{i}\right)$ for $i=1,\cdots,n$. We introduce a time-sharing random variable $Q\sim$ Unif $[1:n]$ independent of $\left(W,X^{n},Y^{n},Z^{n},S^{n}\right)$ and define that $U=\left(Q,U_{Q}\right),V=\left(Q,V_{Q}\right),X=X_{Q},Y=Y_{Q}$, $Z=Z_{Q}$ and $S=S_{Q}$. Using the law of total expectation and the definitions of $Q,S,U,V,Z$ above, we have
@@
$=$ $\mathbb{E}\bigg{[}\mathbb{E}\Big{[}d\Big{(}S_{Q},\hat{s}^{*}\big{(}(U_{Q},V_{Q},Q),Z_{Q}\big{)}\Big{)}\Big{]}\Big{|}Q\bigg{]}=\mathbb{E}\big{[}d\big{(}S,\hat{s}^{*}(U,V,Z)\big{)}\big{]},$ (17)
@@
Note that when the channel to SenRx is degraded with respect to the channel to ComRx, we have the following result.
Theorem 6.
@@
Since $(X,S)-Y-Z$ forms a Markov chain, we have $I\left(U;Y|S\right)\geq I\left(U;Y\right)\geq I\left(U;Z\right)$ holds for all $p(u,x)$. Hence, we obtain that
$I\left(U;Z\right)+I\left(X;Y|U,S\right)\leq I\left(U;Y|S\right)+I\left(X;Y|U,S\right)=I\left(X;Y|S\right),$
and
$I(X;Y,Z|U,V,S)+I(U,V;Z)=I(X;Y|U,V,S)+I(U,V;Z):=I(X;Y|\tilde{U},S)+I(\tilde{U};Z).$
Also, we have $\mathbb{E}\big{[}d\big{(}S,\hat{S}^{*}(U,V,Z)\big{)}\big{]}=\mathbb{E}\big{[}d\big{(}S,\hat{S}^{*}(\tilde{U},Z)\big{)}\big{]}$. Therefore, we get that
$C(D)\leq\max_{P_{\tilde{U}X}}\big{\{}I(X;Y|\tilde{U},S)+I(\tilde{U};Z)\big{|}\mathbb{E}\big{[}d\big{(}S,\hat{S}^{*}(\tilde{U},Z)\big{)}\big{]}\leq D\big{\}}$
according to Theorem 5. Furthermore, we have
$C(D)\geq\max_{P_{UX}}\big{\{}I(X;Y|U,S)+I(U;Z)\big{|}\mathbb{E}\big{[}d\big{(}S,\hat{S}^{*}(U,Z)\big{)}\big{]}\leq D\big{\}}$
according to Theorem 3, which indicates the upper bound 2 coincides with the lower bound in Theorem 3. Thus, we obtain the capacity-distortion function $C(D)$, as shown in (18). ∎
@@
Theorem 6 implies that the optimality of the partial-decoding-based estimation strategy achieved by superposition coding in the degraded case. In addition, the result in Theorem 6 also holds when the Markov chain $X-(Y,S)-Z$ holds.
Referring to the lossy source coding for this ISAC system, we get the following upper bound.
@@
Furthermore, similar to the converse proof in lossy source coding, we have $\sum_{i=1}^{n}I(S_{i};\hat{S}_{i})\leq I(S^{n};\hat{S}^{n})$. And since the Markov chain $S^{n}-Z^{n}-\hat{S}^{n}$ and the composition of auxiliary random variable $U_{i}$, we get
$\sum_{i=1}^{n}I(S_{i};\hat{S}_{i})\leq I(S^{n};\hat{S}^{n})\leq I(S^{n};Z^{n})\leq\sum_{i=1}^{n}I(S_{i},U_{i};Z_{i}),$
which completes the proof. ∎